OpenStudy (anonymous):

(2 + 3i)2 Simplify the expression. A) -5 B) -5 + 12i C) 4 + 9i D) 5 - 12i

RhondaSommer (rhondasommer):

is it \[(2-3i)^2\]

OpenStudy (anonymous):

Yes it is

RhondaSommer (rhondasommer):

2*2+3*3i :)

OpenStudy (anonymous):

thanks

OpenStudy (xmissalycatx):

Wait but that isn't right though.

OpenStudy (anonymous):

You multiply i know it becomes 4 + 9i

OpenStudy (xmissalycatx):

Nonono. That is not correct..

OpenStudy (anonymous):

How? I dont get it

RhondaSommer (rhondasommer):

techincally none of the answers are correct though bc it should be \[i^2\] right @xMissAlyCatx

OpenStudy (poohdessey):

its A

OpenStudy (xmissalycatx):

(2-3i)^2 = (2-3i)(2-3i) = 4 - 12i + 9i^2 i is the square root of -1, so i^2 = -1 So, 4 - 12i + 9i^2 = 4 - 12i - 9 = -5 - 12i

OpenStudy (poohdessey):

the answer is A because 9i becomes 9(-1) and that becomes 4+-9 which equals -5

OpenStudy (xmissalycatx):

So you wouldn't be just simply multiplying, it is like a matter of using the FOIL method first.

OpenStudy (xmissalycatx):

The answer is not A.

RhondaSommer (rhondasommer):

yeah #forgotboutthat

OpenStudy (anonymous):

ah

OpenStudy (xmissalycatx):

Do you understand now?

OpenStudy (poohdessey):

the answer is A

OpenStudy (anonymous):

Kind of,

OpenStudy (xmissalycatx):

@Poohdessey No, it is not. And don't give direct answers even if they are wrong. It is against code of conduct.

OpenStudy (poohdessey):

@xMissAlyCatx whose code of conduct i did explain it. Everyone deserves an explanation, so in the future when he has a problem like it he can do it on his own .... no shade

OpenStudy (xmissalycatx):

The websites code of conduct, want me to link it to you? You didn't explain all you did was give answer.

OpenStudy (poohdessey):

i most definitely did explain it

OpenStudy (xmissalycatx):

But it was incorrect..

OpenStudy (poohdessey):

and you would know because.... right i know im correct lml.... I dont need you to try to tell me what's right or wrong because I'm not the one with the question. sorry, but I am more than capable of working out this math problem on my own.

OpenStudy (xmissalycatx):

It wasn't your math question/post first of all, and I know how to solve it, and I did.. correctly.

OpenStudy (xmissalycatx):

You didn't use the correct method to solve.